OpenStudy (anonymous):

A vendor has 10 balloons for sale: 5 are yellow, 2 are red, and 3 are green. A balloon is selected at random and sold. If the balloon sold is yellow, what is the probability that the next balloon selected at random is also yellow?

OpenStudy (anonymous):

since its 5 yellow would i subtract 1?

Mehek (mehek14):

you had \(\dfrac{5}{10}\) yellow when 1 is sold you have \(\dfrac{4}{9}\) yellow

Mehek (mehek14):

subtract 1 from # of yellow and subtract 1 from total

Mehek (mehek14):

which gives you \(\dfrac{4}{9}\)

OpenStudy (anonymous):

@Mehek14 thanks

Mehek (mehek14):

np